# Parallel AutoRegressive Models for Multi-Agent Combinatorial Optimization

### v2

@@ 4.3 Multi-Agent Encoder
The multi-agent encoder $f_{\theta}$ transforms an input instance $\bm{x}$ into a hidden representation $\bm{h}$. In PARCO, we explicitly model agents and employ separate agent and node embedding layers similar to Son et al. (2024) to project agents and nodes into the same embedding space.
@@
Depending on the problem structure, the initial agent and node embeddings might either be concatenated as $\bm{h}^{(0)}=\text{Concat}(\bm{h}_{a}^{(0)},\bm{h}_{n}^{(0)})$ and passed through $L$ transformer blocks, consisting of multi-head self-attention $\text{MHA}(\bm{h}^{(0)},\bm{h}^{(0)},\bm{h}^{(0)})$ and multi-layer perceptrons (MLPs) as defined in Vaswani et al. (2017). Or, agent and node embeddings are used separately as query and keys/values, respectively, in a cross-attention mechanism $\text{MHA}(\bm{h}_{a}^{(0)},\bm{h}_{n}^{(0)},\bm{h}_{n}^{(0)})$ akin to MatNet (Kwon et al., 2021). The final embeddings $\bm{h}=\{\bm{h}_{a},\bm{h}_{n}\}$ emitted by the last encoder layer contain processed representations of both agents and nodes that capture their interactions as well as the overall problem structure.
@@ 4.4 Communication Layers
At each step $t$ of the decoding process of Equation 3, given the encoded representations $\bm{h}$, we construct dynamic agent queries that capture the current state of both agents and the environment. For each agent $m$, we form a context embedding $\bm{d}_{m}=\text{Concat}(\bm{h}_{a^{m}},\bm{h}_{\delta^{m}_{t}},\bm{h}_{e})$ consisting of the following components: (1) the (static) embedding of the agent $\bm{h}_{a^{m}}$; (2) a projection $\bm{h}_{\delta^{m}_{t}}=\delta^{m}_{t}{\mathbf{W}}_{\delta}\in\mathbb{R}^{d}$ of the agent’s dynamics $\delta^{m}_{t}$ like its current location and capacity (routing) or the time until the agent becomes idle (scheduling); (3) a projection $\bm{h}_{e}=e_{t}{\mathbf{W}}_{\delta}\in\mathbb{R}^{d}$ of the dynamic environment features $e_{t}$ that encode the current problem’s state. These dynamic embeddings are then projected into query vectors $\bm{q}_{m}=\bm{d}_{m}{\mathbf{W}}_{q}$ where ${\mathbf{W}}_{q}\in\mathbb{R}^{3d\times d}$ is a learnable projection matrix.
@@
where Norm denotes a normalization layer (Ioffe & Szegedy, 2015; Zhang & Sennrich, 2019). These layers are inherently agent-count agnostic, which allows PARCO to handle an arbitrary number of agents, making it more flexible and generalizable across different problems. Communication Layers allow agents to coordinate their actions by attending to both the problem structure and other agents’ states while maintaining efficiency through parallel processing.
@@ 4.5 Decoder with Multiple Pointer Mechanism
PARCO’s decoder extends the attention-based pointer mechanism of Kool et al. (2019a) – originally designed for single-agent scenarios and recently applied to sequential multi-agent planning with a single agent at a time (Son et al., 2024; Zheng et al., 2024a) – to handle multiple agents operating in parallel via a Multiple Pointer Mechanism.
@@
$\bm{u}={\beta}\cdot\text{tanh}\left(\frac{\bm{q}(\bm{h}_{n}{\mathbf{W}}^{L}+\xi_{t}{\mathbf{W}}^{L}_{\xi})^{\top}}{\sqrt{d}}\right)$ (7)
with learnable parameters ${\mathbf{W}}^{L}\in\mathbb{R}^{d\times d}$, ${\mathbf{W}}^{L}_{\xi}\in\mathbb{R}^{k_{\xi}\times d}$ and ${\beta}$ is a scale parameter, set to $10$ following Bello et al. (2016) to enhance exploration. The output logits $\bm{u}\in\mathbb{R}^{M\times N}$ are masked by setting infeasible actions given mask $\textbf{M}_{t}$ to minus infinity.The joint probability distribution over all agent actions is then computed as:
@@ 4.6 Conflict Handlers
When sampling from the probability distribution $\bm{p}$ generated by the Multiple Pointer Mechanism, multiple agents may select the same action simultaneously, which can result in an infeasible solution in several CO problems: for instance, in vehicle routing problems, usually only one agent is allowed to visit a customer node. Thus, it is essential to choose how to deal with such a situation effectively. Conflict handling (i.e., tie-breaking) can be achieved by allowing a single agent among a number of agents that are in conflict to continue with its action while others revert to fallback actions, e.g., staying in their current position. A naive approach consists of randomly selecting an agent to perform the new action (Zong et al., 2022). However, this can be suboptimal since it excludes inductive biases that can be leveraged, such as learned representations.
@@
0:  Actions $\bm{a}\in\mathbb{N}^{M}$, Priorities $\bm{p}\in\mathbb{R}^{M}$, Fallback actions $\bm{r}\in\mathbb{R}^{M}$
0:  Resolved Actions $\bm{a}^{\prime}\in\mathbb{N}^{M}$ // Sort indices based on priorities in descending order.
1:  $\sigma\leftarrow\text{argsort}(\bm{p},\text{descending}=\text{True})$ // Reorder actions according to priority.
2:  $\hat{\bm{a}}\leftarrow\bm{a}[\sigma]$ // Initialize conflict mask (all zeros).
3:  $C\leftarrow\bm{0}^{M}$ // Check for conflicts in reordered actions.
4:  for $i=2$ to ${M}$ do
5:     if $\hat{\bm{a}}_{i}\in\{\hat{\bm{a}}_{1},\ldots,\hat{\bm{a}}_{i-1}\}$ then
6:        $C_{i}\leftarrow 1$
7:     end if
8:  end for// Resolve conflicts by assigning fallback actions.
9:  $\hat{\bm{a}}\leftarrow(1-C)\odot\hat{\bm{a}}+C\odot\bm{r}$ // Reorder resolved actions back to original order.
10:  $\bm{a}^{\prime}\leftarrow\hat{\bm{a}}[\sigma^{-1}]$
In PARCO, we propose Priority-based Conflict Handlers that leverage priorities as a tie-breaking rule. Such priorities can be based on heuristics – such as giving priority to agents close to completion or whose action results in the smallest immediate cost – or on learned priorities. In the latter case, the probability values of the selected actions $p(\bm{a}_{t})$ serve as an indicator for prioritizing certain agents: the higher its value, the higher the priority learned by the model to have that agent win the tie-break. Algorithm 1 shows an efficient vectorized implementation of the Priority-based Conflict Handler algorithm. In practice, we augment the conflict handler $\psi$ from Section 4.2 with priorities $\bm{p}\coloneq p(\bm{a}_{t})$ and fallback actions $\bm{r}$, i.e. $\psi(\cdot)\coloneq\psi(\bm{a}_{t},\bm{p},\bm{r})$. Fallback actions in PARCO correspond to “do nothing” operations – maintaining an agent’s current position (routing) or keeping a machine idle (scheduling) – which may result in slightly more solution construction steps but do not affect the final solution $\bm{a}$. This approach effectively handles conflicts by allowing the affected agents to reconsider their choices given the actions of (preceding) agents in the next decoding step.
@@ 4.7 Training Scheme
PARCO is a centralized multi-agent decision-making framework and can thus be trained by RL algorithms proposed in the single-agent NCO literature. We train PARCO by employing the REINFORCE gradient estimator (Williams, 1992) with a shared baseline as outlined by Kwon et al. (2020) and Kim et al. (2022):
@@ 5 Experiments
We assess the effectiveness of PARCO on representative multi-agent CO problems, spanning both routing and scheduling domains. Specifically, we evaluate its performance on two challenging routing problems – the min-max heterogeneous capacitated vehicle routing problem (HCVRP) and the open multi-depot capacitated pickup and delivery problem (OMDCPDP) – as well as a scheduling problem, the flexible flow shop problem (FFSP).
@@ 5.1 Experimental Settings
We experiment on a workstation equipped with 2 Intel(R) Xeon(R) Gold 6338 CPUs and $8$ NVIDIA RTX 4090 graphic cards with $24$ GB of VRAM each. Training runs of PARCO take less than 24 hours each. During inference, we employ only one CPU and a single GPU. Additional details are available throughout Appendix B.
@@
Problem. The min-max HCVRP involves ${M}$ agents serving customer demands while adhering to heterogeneous vehicle capacity constraints. Each vehicle can replenish its load by returning to the depot. The objective is to minimize the longest route taken by any agent (min-max), ensuring balanced workload distribution. Traditional solvers. We include state-of-the-art SISRs (Christiaens & Vanden Berghe, 2020), Genetic Algorithm (GA) (Karakatič & Podgorelec, 2015) and Simulated Annealing (SA) (İlhan, 2021). Neural baselines. We evaluate the sequential planning baselines Attention Model (AM) (Kool et al., 2019a), Equity Transformer (ET) (Son et al., 2024) and Decoupling Partition and Navigation (DPN) (Zheng et al., 2024a), and autoregressive models with agent selection DRLLi (Li et al., 2022) and state-of-the-art learning method 2D-Ptr (Liu et al., 2024b). Additional problem and experimental details are available in Section A.1 and Section B.1, respectively.
@@
Problem. The OMDCPDP is a challenging problem arising in last-mile delivery settings where $M$ agents starting from different locations (i.e., multiple depots) must pick up and deliver parcels without returning to their starting point (i.e., open). Agents have a capacity constraint for orders that can be carried out as a stacking limit: a tour can include more pickups than the constraint, but the agent must deliver corresponding orders so that carrying capacity is freed. The goal is to minimize the lateness, i.e., the sum of delivery arrival times. Traditional solvers. We include the popular and efficient optimization suite Google OR-Tools (Furnon & Perron, 2024) as a classical baseline. Neural baselines. We evaluate models specializing in pickup and delivery problems, including the autoregressive Heterogeneous Attention Model (HAM) (Li et al., 2021a) for sequential planning and MAPDP (Zong et al., 2022) for parallel planning. Additional problem and experimental details are available in Section A.2 and Section B.2, respectively.
@@
Problem. In FFSP, $N$ jobs must be processed by $M$ machines divided equally in $S$ stages. Jobs follow a specified sequence through these stages. Within each, any available machine can process the job, with the key constraint that no machine can handle multiple jobs simultaneously. The goal is to schedule the jobs so that all jobs are finished in the shortest time possible. Traditional solvers. We incorporate the widely used and powerful Gurobi solver (Gurobi Optimization, LLC, 2024) as a baseline. Furthermore, we include dispatching rules Random and Shortest Job First (SJF), Particle Swarm Optimization (PSO) (Singh & Mahapatra, 2012) and Genetic Algorithm (GA) (Hejazi & Saghafian, 2005). Neural baselines. Notable benchmarks include the Matrix Encoding Network (MatNet) (Kwon et al., 2021) which demonstrates superior performance on FFSP. Additional problem and experimental details are available in Section A.3 and Section B.3, respectively.
Table 1: Main results with different configurations in terms of problem size $N$ and number of agents $M$. The lower the better ($\downarrow$).
@@
Figure 3: Analysis on PARCO components. (a) Full Communication Layers provide the best performance. (b) Average number of decoding steps required to construct solutions (bar heights) and training times (lines) for different models. (c) Random and Priority-based Conflict Handlers variants with heuristics and learned priorities with learned providing the best performance.
@@ 5.2 Experimental Results
Table 2: Large-scale generalization for unseen numbers of nodes $N$ and agents $M$ (up to 10×10$\times$10 × of those seen during training).
$N$ 500 1000
@@
We present the main empirical results for HCVRP, OMDCPDP, and FFSP in Table 1, reporting the average objective function values (Obj.), gaps to the best-known solutions, and inference times for solving each single instance of each problem. For neural baselines, we evaluate both greedy (g.) and sampling (s.) performance, using 1280 sampled solutions for routing problems and 128 for FFSP.
In HCVRP, PARCO outperforms all neural baselines in solution quality and speed while providing solutions at a fraction of the solving time required by traditional solvers. In OMDCPDP, our model surpasses all baselines, including OR-Tools. Notably, while the AR baseline HAM struggles with a larger number of agents $M$, PARCO’s parallel AR method maintains strong performance across all scales. In FFSP, PARCO outperforms traditional solvers (e.g., Gurobi cannot find solutions in time for $N>20$), dispatching rules, and MatNet in all tested scenarios while being more than 4×4$\times$4 × faster. Furthermore, similar to our results in routing, PARCO’s advantage becomes even more pronounced in instances with a larger number of agents where PARCO generates higher-quality schedules through effective agent coordination at a fraction of the cost of MatNet.
5.3 Analysis
Effect of Communication Layers
We showcase the importance of Communication Layers in Figure 3(a). We benchmark different ways to modify decoder queries (see Section 4.5) with 1) No communication (only context features), 2) MLP, 3) MHA, 4) our transformer-based Communication Layers. Our Communication Layers consistently outperform other methods. Figure 3(b) further shows the effect of communication, which can not only improve performance but also reduce training time by reducing the number of steps, thereby offsetting the additional latency introduced by the Communication Layers.
@@ 5.3 Analysis
We study the effect of conflict handlers in Figure 3(c). We compare 1) the random handler from MAPDP and our proposed Priority-based Conflict Handlers in different configurations, namely with priorities based on heuristics, including 2) “smallest” prioritizing the agent with the lowest cost so far, and 3) “closest”, prioritizing agents closer to the corresponding node, and finally 4) “learned” with learned priorities (see Section 4.6). PARCO’s Priority-based Conflict Handlers with learned priorities consistently outperform other methods, which also enjoy a relative reduction in the number of steps for constructing a solution, e.g., with a $4\%$ reduction in conflict rates.
@@
We study the zero-shot large-scale generalization performance of PARCO on OMDCPDP and report the results in Table 2 for out-of-distribution numbers of nodes $N$ and agents $M$, both up to 10×10$\times$10 × those seen in training. The autoregressive HAM baseline cannot generalize well to such scales due to the lack of communication and robust parallel construction. We note that MAPDP cannot be applied to an unseen number of agents because of its inflexible decoder structure. Conversely, our method outperforms all baselines, even including OR-Tools with a 10-minute solving time per instance for $N=1000$, making PARCO a powerful candidate for real-time deployment.
Figure 4: PARCO vs AR inference time. PARCO can construct solutions faster with more agents $M$ thanks to parallel decoding.
@@
Finally, we showcase PARCO’s scalability against autoregressive methods in Figure 4. Notably, compared to AR models (e.g., DPN, HAM), PARCO achieves significantly faster solution speeds, with efficiency improving as the number of agents $M$ increases. Thanks to parallel decoding, fewer solution construction steps for larger $M$ lead to substantially lower latency.

### v3

@@ 4.3 Multi-Agent Encoder
The multi-agent encoder $f_{\theta}$ transforms an input instance $\bm{x}$ into a hidden representation $\bm{h}$. In PARCO, we explicitly model agents and employ separate agent and node embedding layers similar to Son et al. [84] to project agents and nodes into the same embedding space.
@@
Depending on the problem structure, the initial agent and node embeddings might either be concatenated as $\bm{h}^{(0)}=\text{Concat}(\bm{h}_{a}^{(0)},\bm{h}_{n}^{(0)})$ and passed through $L$ transformer blocks, consisting of multi-head self-attention $\text{MHA}(\bm{h}^{(0)},\bm{h}^{(0)},\bm{h}^{(0)})$ and multi-layer perceptrons (MLPs) as defined in Vaswani et al. [88]. Or, agent and node embeddings are used separately as query and keys/values, respectively, in a cross-attention mechanism $\text{MHA}(\bm{h}_{a}^{(0)},\bm{h}_{n}^{(0)},\bm{h}_{n}^{(0)})$ akin to MatNet [50]. The final embeddings $\bm{h}=\{\bm{h}_{a},\bm{h}_{n}\}$ emitted by the last encoder layer contain processed representations of both agents and nodes that capture their interactions as well as the overall problem structure.
@@ 4.4 Communication Layers
At each step $t$ of the decoding process of Eq.˜2, given the encoded representations $\bm{h}$, we construct dynamic agent queries that capture the current state of both agents and the environment. For each agent $m$, we form a context embedding $\bm{d}_{m}=\text{Concat}(\bm{h}_{a^{m}},\bm{h}_{\delta^{m}_{t}},\bm{h}_{e})$ consisting of the following components: (1) the (static) embedding of the agent $\bm{h}_{a^{m}}$; (2) a projection $\bm{h}_{\delta^{m}_{t}}=\delta^{m}_{t}{\mathbf{W}}_{\delta}\in\mathbb{R}^{d}$ of the agent’s dynamics $\delta^{m}_{t}$ like its current location and capacity (routing) or the time until the agent becomes idle (scheduling); (3) a projection $\bm{h}_{e}=e_{t}{\mathbf{W}}_{\delta}\in\mathbb{R}^{d}$ of the dynamic environment features $e_{t}$ that encode the current problem’s state. These dynamic embeddings are then projected into query vectors $\bm{q}_{m}=\bm{d}_{m}{\mathbf{W}}_{q}$ where ${\mathbf{W}}_{q}\in\mathbb{R}^{3d\times d}$ is a learnable projection matrix.
@@
where Norm denotes a normalization layer [33, 100]. These layers are inherently agent-count agnostic, allowing PARCO to handle arbitrary numbers of agents, making it more flexible and generalizable across different problems. Communication Layers allow agents to coordinate their actions by attending to both the problem structure and other agents’ states while maintaining efficiency through parallel processing.
@@ 4.5 Decoder with Multiple Pointer Mechanism
PARCO’s decoder improves the AR pointer mechanism [89, 47] – originally designed for single-agent scenarios and recently applied to sequential multi-agent planning with a single agent at a time [84, 104] – to handle multiple agents operating in parallel via a Multiple Pointer Mechanism.
@@
$\bm{u}={\beta}\cdot\text{tanh}\left(\frac{\bm{q}^{\prime}(\bm{h}_{n}{\mathbf{W}}^{L}+\xi_{t}{\mathbf{W}}^{L}_{\xi})^{\top}}{\sqrt{d}}\right)$ (6)
with learnable parameters ${\mathbf{W}}^{L}\in\mathbb{R}^{d\times d}$, ${\mathbf{W}}^{L}_{\xi}\in\mathbb{R}^{k_{\xi}\times d}$ and $\beta$} is a scale parameter, set to $10$ following Bello et al. [3] to enhance exploration. The output logits $\bm{u}\in\mathbb{R}^{M\times N}$ are masked by setting infeasible actions given mask $\textbf{M}_{t}$ to -$\infty$. The joint probability distribution over all agent actions becomes:
@@ 4.6 Conflict Handlers
When sampling from the probability distribution $\bm{p}$ generated by the Multiple Pointer Mechanism, multiple agents may select the same action simultaneously, which can result in an infeasible solution in several CO problems – for instance, in vehicle routing problems, usually only one agent is allowed to visit a customer node – and it becomes essential how to deal with such a situation effectively. Conflict handling (i.e., tie-breaking) can be achieved by allowing a single agent among a number of agents that are in conflict to continue with its action while others revert to fallback actions, e.g., staying in their current position. A simple approach introduced by Zong et al. [108] consists of randomly selecting an agent to perform the new action. However, this can be suboptimal since it excludes inductive biases that can be leveraged, such as learned representations.
@@
1:Actions $\bm{a}\in\mathbb{N}^{M}$, Priorities $\bm{p}\in\mathbb{R}^{M}$, Fallback actions $\bm{r}\in\mathbb{R}^{M}$
2:Resolved Actions $\bm{a}^{\prime}\in\mathbb{N}^{M}$
[…]
6:for $i=2$ to M{M} do // Check for conflicts in reordered actions
7:  if $\hat{\bm{a}}_{i}\in\{\hat{\bm{a}}_{1},\ldots,\hat{\bm{a}}_{i-1}\}$ then
8:   $C_{i}\leftarrow 1$ // $C_{i}=1$ indicates a conflict for index $i$
9:  end if
10:end for
[…]
In PARCO, we propose Priority-based Conflict Handlers that leverage priorities as a tie-breaking rule. Such priorities can be based on heuristics – such as giving priority to agents close to completion or whose action results in the smallest immediate cost – or on learned priorities. In the latter case, the model output probability values of the selected actions $p(\bm{a}_{t})$ serve as an indicator for prioritizing certain agents: the higher their value, the higher the priority learned by the model to have those agents win the tie-break.
Algorithm˜1 shows our efficient vectorized implementation of the Priority-based Conflict Handler algorithm. In practice, we augment the conflict handler $\psi$ from §˜4.2 with priorities $\bm{p}\coloneq p(\bm{a}_{t})$ and fallback actions $\bm{r}$, i.e. $\psi(\cdot)\coloneq\psi(\bm{a}_{t},\bm{p},\bm{r})$. Fallback actions in PARCO correspond to “do nothing” operations – maintaining an agent’s current position (routing) or keeping a machine idle (scheduling) – which may result in slightly more solution construction steps but do not affect the final solution $\bm{a}$. This approach effectively handles conflicts by allowing the affected agents to reconsider their choices given the actions of (preceding) agents in the next decoding step.
@@ 4.7 Training Scheme
PARCO is a centralized multi-agent decision-making framework and can thus be trained by RL algorithms proposed in the single-agent NCO literature. We train PARCO by employing the REINFORCE gradient estimator [92] with a shared baseline as outlined by Kwon et al. [49] and Kim et al. [43]:
@@ 5 Experiments
We assess the effectiveness of PARCO on representative multi-agent CO problems, spanning both routing and scheduling domains. Specifically, we evaluate its performance on two challenging routing problems – the min-max heterogeneous capacitated vehicle routing problem (HCVRP) and the open multi-depot capacitated pickup and delivery problem (OMDCPDP) – as well as a scheduling problem, the flexible flow shop problem (FFSP). Experimental details are available through Appendix˜B [^1] Source code is available at https://github.com/ai4co/parco.
@@ 5.1 Experimental Settings
Problem. The min-max HCVRP involves M{M} agents serving customer demands while adhering to heterogeneous vehicle capacity constraints. Each vehicle can replenish its load by returning to the depot. The objective is to minimize the longest route taken by any agent (min-max), ensuring balanced workload distribution. Traditional solvers. We include state-of-the-art SISRs [12], Genetic Algorithm (GA) [39] and Simulated Annealing (SA) [32]. Neural baselines. We evaluate the sequential planning baselines Attention Model (AM) [47], Equity Transformer (ET) [84] and Decoupling Partition and Navigation (DPN) [104], and autoregressive models with agent selection DRLLi [55] and state-of-the-art learning method 2D-Ptr [63]. Additional problem and experimental details are available in §˜A.1 and §˜B.1, respectively.
@@
Problem. The OMDCPDP is a challenging problem arising in last-mile delivery settings where $M$ agents starting from different locations (i.e., multiple depots) must pick up and deliver parcels without returning to their starting point (i.e., open). Agents have a capacity constraint for orders that can be carried out as a stacking limit: a tour can include more pickups than the constraint, but the agent must deliver corresponding orders so that carrying capacity is freed. The goal is to minimize the lateness, i.e., the sum of delivery arrival times. Traditional solvers. We include the popular and efficient optimization suite Google OR-Tools [19] as a classical baseline. Neural baselines. We evaluate models specializing in pickup and delivery problems, including the autoregressive Heterogeneous Attention Model (HAM) [54] for sequential planning and MAPDP [108] for parallel planning. Additional problem and experimental details are available in §˜A.2 and §˜B.2, respectively.
@@
Problem. In FFSP, $N$ jobs must be processed by $M$ machines divided equally in $S$ stages. Jobs follow a specified sequence through these stages. Within each, any available machine can process the job, with the key constraint that no machine can handle multiple jobs simultaneously. The goal is to schedule the jobs so that all jobs are finished in the shortest time possible. Traditional solvers. We incorporate the widely used and powerful Gurobi solver [24] as a baseline. Furthermore, we include dispatching rules Random and Shortest Job First (SJF), Particle Swarm Optimization (PSO) [81] and Genetic Algorithm (GA) [25]. Neural baselines. Notable benchmarks include the Matrix Encoding Network (MatNet) [50] which demonstrates superior performance on FFSP. Additional problem and experimental details are available in §˜A.3 and §˜B.3, respectively.
Table 1: Main results on different problems with different configurations for problem size $N$ and number of agents $M$. For all metrics, the lower the better ($\downarrow$).
@@ 5.2 Experimental Results
We report the main empirical results for HCVRP, OMDCPDP, and FFSP in Table˜1, with average objective function values (Obj.), gaps to the best-known solutions, and inference times for solving each single problem instance. For neural baselines, we evaluate both greedy (g.) and sampling (s.) performance, using 1280 sampled solutions for routing problems and 128 for FFSP.
In HCVRP, PARCO outperforms all neural baselines in solution quality and speed while providing solutions at a fraction of the solving time required by traditional solvers. In OMDCPDP, our model surpasses all baselines, including OR-Tools. Notably, while the AR baseline HAM struggles with a larger number of agents $M$, PARCO’s parallel AR method maintains strong performance across all scales. In FFSP, PARCO outperforms traditional solvers (e.g., Gurobi cannot find solutions in time for $N>20$), dispatching rules, and MatNet in all tested scenarios while being more than $4\times$ faster. Furthermore, similar to our results in routing, PARCO’s advantage becomes even more pronounced in instances with a larger number of agents where PARCO generates higher-quality schedules through effective agent coordination at a fraction of the cost of MatNet.
5.3 Analysis
Effect of Communication Layers
We showcase the importance of Communication Layers in Fig.˜3(a). We benchmark different ways to obtain decoder queries (see §˜4.5) with 1) No communication (W/o Comm., i.e., with only context features), 2) MLP, 3) MHA, 4) our transformer-based Communication Layers (Comm.). Our Communication Layers consistently outperform other methods. Fig.˜3(b) shows decoding steps and training times on the FFSP for $N=50$. PARCO greatly reduces the number of steps and training times, with Communication Layers further reducing them through better coordination, especially at a higher number of agents $M$.
Table 2: Generalization for unseen numbers of nodes $N$ and agents $M$ (up to $10\times$ those seen during training).
NN 500 1000
@@ 5.3 Analysis
(a) Effect of communication
(b) Decoding steps and training time
(c) Effect of conflict handlers
Figure 3: Analysis of PARCO components.
@@
In Fig.˜3(c), we compare 1) the random handler from MAPDP and our proposed Priority-based Conflict Handlers in different configurations, namely with priorities based on simple heuristics as 2) “smallest” prioritizing the agent with the lowest cost so far, and 3) “closest”, prioritizing agents closer to the corresponding node, and finally 4) “learned” based on model output probabilities. The latter consistently outperforms other methods, which also enjoy a relative reduction in the number of steps for constructing a solution, e.g., with a $4\%$ reduction in conflict rates.
@@
We study the zero-shot large-scale generalization performance of PARCO in the OMDCPDP and report the results in Table˜2 for out-of-distribution numbers of nodes $N$ and agents $M$, both up to $10\times$ those seen in training. We find that the AR HAM baseline cannot generalize well to such scales due to the lack of communication and robust parallel construction, while MAPDP cannot be applied to an unseen $M$ because of its inflexible decoder structure.
Figure 4: PARCO vs AR inference time. PARCO constructs solutions faster with more agents $M$.
Conversely, our method outperforms all baselines, including OR-Tools with a 10-minute solving time per instance for $N=1000$, making PARCO a strong candidate for real-time deployment.
@@
Finally, we showcase PARCO’s speedups against autoregressive methods in Fig.˜4. Notably, compared to AR models (e.g., DPN, HAM), PARCO achieves significant speedups of $3.3\times$ up to $24.7\times$, with inference time decreasing as the number of agents $M$ increases: thanks to parallel decoding, fewer solution construction steps for larger $M$ lead to substantially lower latency.